OpenStudy (anonymous):

How do I find the zeros of the equation given that one of the zeros is x = 2 ? x^3 - 7x + 6 = 0

OpenStudy (anonymous):

Since one of the zeroes is 2, you know that you can factor the given function so that it looks like this: \[\text{If }p(x)=x^3-7x+6, \text{ then }p(x)=(x-2)q(x).\]

OpenStudy (anonymous):

To find \(q(x)\), try dividing \(p(x)\) by \(x-2\), either by long division or synthetic division.

OpenStudy (anonymous):

q(x) is quotient? synthetic division gives me \[x^3 + 2x^2 -3x\]

OpenStudy (anonymous):

that means my equation is \[p(x) = (x-2)(x^3+2x^2-3x)\] or am I completely off?

OpenStudy (anonymous):

Your result from synthetic division is a bit off, yeah. The way you have \(p(x)\), if you were to distribute the \(x-2\), you'd get another polynomial of degree higher than the given polynomial, which can't be true.

OpenStudy (anonymous):

Your only mistake there was that you should have gotten \(q(x)=x^2+2x-3\). So, you have \(x^3-7x+6=(x=2)(x^2+2x-3)\).

OpenStudy (anonymous):

 where's my mistake? (I don't want to make it again)